The cost structure of a manufacturer of microchips is described in the table shown below.

The​ firm's fixed costs equal ​$25​,000 per day. Calculate the average variable cost​, average fixed cost​, and average total cost at each output level.

​(Your answers should be rounded to the nearest cent.​)

output (microchips per day) total cost of the output average variable cost average fixed cost average total cost
0 25,000 - - -
25,000 75,000
50,000 110,000
75,000 165,000
100,000 235,000
125,000 340,000
150,000 480,000
Answer #1

Answer:

Variable Cost (VC)= Total Cost (TC) - Fixed Cost((FC)

Average Variable Cost(AVC) = AVC / Q

Average Fixed Cost(AFC) = FC / Q

Average Total Cost(ATC) = TC / Q

Output(Q)

Total Cost(TC)

Fixed Cost(FC)

Variable Cost(VC) = TC - FC

Average Variable Cost(AVC) = AVC / Q

Average Fixed Cost(AFC) = FC / Q

Average Total Cost(ATC) = TC / Q

0

25,000

25,000

0

-

-

-

25,000

75,000

25,000

50,000

2

1

3

50,000

110,000

25,000

85,000

1.7

0.5

2.2

75,000

165,000

25,000

140,000

1.87

0.33

2.2

100,000

235,000

25,000

210,000

2.1

0.25

2.35

125,000

340,000

25,000

315,000

2.52

0.2

2.72

150,000

480,000

25,000

455,000

3.03

0.17

3.2